OpenStudy (anonymous):

If sin Θ = -√3/2 and π < Θ < 3π/2 , what are the values of cos Θ and tan Θ?

OpenStudy (solomonzelman):

 \(\LARGE\color{blue}{ \bf sin(60)=\frac{\sqrt{3} }{ 2} }\) Hence, \(\LARGE\color{blue}{ \bf sin(-60)=\frac{-\sqrt{3} }{ 2} }\)

OpenStudy (anonymous):

I cant see that problem clearly it has this \(\LARGE\color{blue}{ \bf sin(60)=\frac{\sqrt{3} }{ 2} }\) Hence, \(\LARGE\color{blue}{ \bf sin(-60)=\frac{-\sqrt{3} }{ 2} }\)